# AutoProSAM: Automated Prompting SAM for 3D Multi-Organ Segmentation

### v3

@@ Abstract
Segment Anything Model (SAM) is one of the pioneering prompt-based foundation models for image segmentation and has been rapidly adopted for various medical imaging applications. However, in clinical settings, creating effective prompts is notably challenging and time-consuming, requiring the expertise of domain specialists such as physicians. This requirement significantly diminishes SAM’s primary advantage—its interactive capability with end users—in medical applications. Moreover, recent studies have indicated that SAM, originally designed for 2D natural images, performs suboptimally on 3D medical image segmentation tasks. This subpar performance is attributed to the domain gaps between natural and medical images and the disparities in spatial arrangements between 2D and 3D images, particularly in multi-organ segmentation applications. To overcome these challenges, we present a novel technique termed AutoProSAM. This method automates 3D multi-organ CT-based segmentation by leveraging SAM’s foundational model capabilities without relying on domain experts for prompts. The approach utilizes parameter-efficient adaptation techniques to adapt SAM for 3D medical imagery and incorporates an effective automatic prompt learning paradigm specific to this domain. By eliminating the need for manual prompts, it enhances SAM’s capabilities for 3D medical image segmentation and achieves state-of-the-art (SOTA) performance in CT-based multi-organ segmentation tasks.
@@ 1 Introduction
Despite their extensive training, large foundation models such as SAM struggle to adapt to medical images. This is primarily due to the inherent differences between medical and natural images and the challenges associated with effective prompting (as shown in Fig. 1). Recent efforts have attempted to extend the success of SAM to medical image segmentation tasks [43, 39, 21, 7, 33]. However, the demonstrated performance has exhibited reduced precision and stability, particularly in more intricate segmentation tasks in 3D medical images characterized by smaller sizes, irregular shapes, and lower contrast properties compared to 2D natural images [7]. Fig. 1(A) illustrates this oversight of substantial domain disparities between natural and medical images. Furthermore, unlike natural images, medical images in various formats are not easily interpretable by general users. The prompts, intended to guide the foundation models to focus on a particular location, become a performance bottleneck as users struggle to identify the region of interest. Therefore, generating effective prompts for these models necessitates the involvement of domain experts and physicians. This reliance on labor-intensive manually generated prompts [5, 33] hampers its successful application, particularly in multi-organ medical image segmentation tasks, as demonstrated in Fig. 1(B). As a result, these foundation models have been shown to underperform in medical image segmentation tasks compared to state-of-the-art (SOTA) models specifically designed for medical imagery [7].
@@
  • We conduct extensive experiments and analysis on three public and one private multi-organ segmentation datasets, demonstrating that the proposed AutoProSAM achieves superior performance in medical image segmentation tasks compared to the SOTAs.
@@ 2.2 Parameter-efficient Model Fine-Tuning
Given the extensive utilization of foundation models, the concept of parameter-efficient fine-tuning has garnered significant attention. Existing methods for efficient fine-tuning can be categorized into three groups [3]. First, addition-based methods incorporate lightweight adapters [27, 37] or prompts [20, 13] into the original model, focusing solely on adjusting these parameters. Second, specification-based methods [40, 8] concentrate on selecting a small subset of the original parameters for tuning. Third, reparameterization-based methods [10] leverage low-rank matrices to approximate parameter updates. Recently, some researchers have extended pre-trained image models to encompass video comprehension [27] or volumetric segmentation [37]. However, these methods treat the additional dimension as a ”word group” and employ specialized modules to aggregate information along the word dimension. In contrast, our work considers all three dimensions simultaneously and directly adapts the trained transformer block to capture patterns in 3D medical image inputs.
@@ 3 Method
In this section, we explain how to modify the original SAM architecture, initially developed for 2D natural images, to work with 3D volumetric medical images for segmentation tasks. We begin by offering a brief overview of the SAM framework (as depicted in Fig. 2), followed by a detailed explanation of the adjustments made to the image encoder, prompt encoder, and mask decoder.
@@ 3.2 Handling 3D Medical Inputs
To enhance SAM’s capability in processing 3D medical images, we propose an adaptation strategy named AutoProSAM, as illustrated in Fig. 2A. This strategy has two primary objectives: firstly, to enable the model to directly learn 3D spatial patterns, and secondly, to ensure continuity by inheriting most parameters from the pre-trained model while introducing easily adjustable incremental parameters. The detailed design is elaborated as follows.
@@ 3.2.4 Bottleneck Modifications
By making the above adjustments, we can smoothly transition the 2D ViT into a 3D ViT, reusing most parameters. However, fully fine-tuning the 3D ViT can be resource-intensive. To address this, we propose using a lightweight adapter approach for efficient fine-tuning. The adapter comprises a down-projection linear layer and an up-projection linear layer, represented as $\operatorname{Adapter}(\mathbf{X})=\mathbf{X}+\operatorname{Act}(\mathbf{X}W_{\mathrm{Down}})W_{\mathrm{Up}}$. Here, $\mathbf{X}\in\mathbb{R}^{N\times C}$ is the original feature representation, $W_{\mathrm{Down}}\in\mathbb{R}^{C\times N^{\prime}}$ and $W_{\mathrm{Up}}\in\mathbb{R}^{N^{\prime}\times C}$ are down-projection and up-projection layers, and $\mathrm{Act}(\cdot)$ is the activation function. To enhance 3D spatial awareness, we include a depth-wise 3D convolution after the down-projection layer, as shown in Fig. 2B. This enhancement improves the adapter’s utilization of 3D spatial cues.
@@ 3.3 Auto Prompt Generator
To overcome these limitations, we propose to use an Auto Prompt Generator instead of positional encoding to represent the prompt. The whole process is illustrated in Fig. 2C. Instead of using manually generated points or bounding boxes, we directly take the output feature map after the last block of attention and depth adapter operation. This APG follows a fully convolutional neural (FCN) based encoder-decoder design that resembles 3D UNet [32]. This generator boasts a lightweight structure, leveraging 3D-based convolution operations, and can be effortlessly learned from scratch. This enables precise prompt generation tailored to different medical segmentation tasks. Notably, it eliminates the need for additional manually generated prompts, simplifying and expediting the multi-class medical image segmentation tasks.
@@ 3.4 Mask Decoder
The mask decoder in SAM was originally designed to be lightweight, primarily using stacks of 2D convolution layers. In our updated version, we have replaced these 2D convolutions with 3D convolutions to enable direct 3D mask generation. The initial design of the decoder, which lacks progressive upsampling and skip connections, is effective for natural images that typically feature large objects with distinct boundaries. However, in the context of volumetric medical image segmentation, it is well-recognized that U-shaped networks with skip connections at multiple levels are crucial for success [11, 34]. This is due to the smaller size of objects in medical images and their often indistinct boundaries, necessitating networks that can preserve higher-resolution details for better discrimination. Therefore, the use of U-shaped architectures with skip connections becomes imperative.
To address this need while keeping the design lightweight, we incorporate a multi-layer aggregation mechanism (MLAM) [44, 7] in our decoder. We utilize the intermediate feature maps from stages 1-4 (as depicted in Fig. 2A) of the image encoder, along with the prompt embedding from the APG, to enrich the mask feature map without compromising efficiency. For improved resolution detail, we upsample the mask feature map to match the original resolution, then concatenate it with the original image. This concatenated map is fused using another 3D convolution to generate the final mask. This method effectively combines high-resolution details with the original image data in the mask generation process. We have streamlined the original SAM by focusing solely on a targeted downstream task and omitting features like multi-task generation and ambiguity awareness. The backbone of the mask decoder primarily comprises lightweight 3D convolutional layers, which are known for their ease of optimization. Consequently, we train all decoder parameters from scratch.
@@ 4.1 Datasets
AMOS: We actually use the CT parts in the original AMOS dataset [12], not including their MRI parts. The publicly accessible AMOS-CT dataset [12] consists of 200 multi-contrast abdominal CT scans for training and 100 for testing. These scans are annotated for 15 anatomies, including spleen, right kidney, left kidney, gallbladder, esophagus, liver, stomach, aorta, inferior vena cava, pancreas, right adrenal gland, left adrenal gland, duodenum, bladder, and prostate/uterus.
@@ 4.2 Implementation Details
We provide detailed pre-processing information about the interpolation to a fixed voxel spacing, HU ranges, and intensity normalization in Tab. 1.
@@ 4.3 Loss Function
For training the AutoProSAM (as shown in Fig. 2A), a combination of Dice loss and Cross-Entropy loss is used to assess the alignment between the predicted mask and the ground truth on a pixel-wise basis. The objective function for the segmentation head is defined as follows:
@@ 5 Results & Discussion
Model Tuned Params. BTCV AMOS CT-ORG Institutional Pelvic
(M) mDice $\uparrow$ mNSD $\uparrow$ mDice $\uparrow$ mNSD $\uparrow$ mDice $\uparrow$ mNSD $\uparrow$ mDice $\uparrow$ mNSD $\uparrow$
nnUNet [11] 31.18 84.34 73.21 87.43 77.12 85.51 75.60 88.04 83.87
AttUNet [36] 40.79 83.67 74.35 83.75 72.23 85.13 74.64 88.06 84.12
UNet++ [47] 28.45 79.33 72.84 84.32 73.34 84.87 74.21 87.53 82.65
nnFormer [46] 150.14 83.51 71.65 84.52 70.06 82.75 70.31 87.60 82.13
UNETR [9] 93.02 85.47 74.35 77.24 60.58 83.13 71.56 86.68 81.54
SwinUNETR [34] 62.83 86.58 75.26 86.19 74.83 84.34 72.32 89.12 83.54
Ours 26.53 87.15 78.83* 88.65* 79.41* 85.12 76.37* 91.30* 84.35*
@@ 5.1 Comparison with SOTAs
We extensively compare our model with the SOTA 3D medical image segmentation approaches, including the most recent Transformer-based methods including UNETR [9], SwinUNETR [34], and nnFormer [46], as well as CNN-based methods such as nnUNet [11], AttUNet [36], and UNet++ [47]. As indicated in Tab. 2, we observe that the proposed AutoProSAM generally outperforms other SOTA methods in both Dice and NSD metrics across all four datasets.
@@
The superiority of our AutoProSAM model is further highlighted by qualitative results. As depicted in Fig. 3, qualitative comparisons of the predicted masks from various segmentation models reveal that our AutoProSAM yields visually superior mask predictions, particularly in terms of more accurate boundary delineation, when compared to its SOTA counterparts. In summary, while the original SAM displayed relatively weaker performance in medical image segmentation tasks as compared to SOTA methods [22], our adapted design registers significant enhancements and generalization to 3D medical image segmentation.
We also randomly selected samples from the CT-ORG dataset for more visual comparisons of different methods. As depicted in Fig. 4, all methods generally achieve precise prediction masks for organs that are relatively easy to segment, such as lungs and bones. However, our AutoProSAM also provides more consistently accurate predictions than other baselines for organs like the liver, kidney, and bladder.
@@ 5.2 Comparison with SAM-based Methods
We further compare our method with existing SAM-based methods for multi-class medical image segmentation, including the original SAM and MedSAM [21]. MedSAM can be implemented with either full fine-tuning or partial fine-tuning, where only the prompt encoder and mask decoder are updated using point prompts. Other adaptation methods were not considered, as they either require collecting a large number of medical images for fine-tuning (e.g., Medical-SAM-Adapter [39], SAM-Med2D [1]), or they do not specifically target multi-organ segmentation tasks (e.g., 3DSAM-Adapter [7]).
@@
The outcomes are meticulously outlined in Tab. 3, underscoring how our adaptation strategy surpasses all existing methods. Notably, our approach outshines the second-best technique by a margin of 3% in terms of BTCV segmentation Dice, 5% for NSD, and 6% concerning AMOS segmentation Dice, accompanied by a substantial 9% for NSD. Impressively, it even outperforms the complete fine-tuning variant of MedSAM, even when considering parameter-efficient fine-tuning. These results effectively validate our hypothesis that parameters pre-trained on 2D images can be effectively harnessed to grasp 3D spatial features with only minor adjustments. Moreover, our approach of treating all dimensions equivalently emerges as a superior strategy compared to interpreting the depth dimension as a distinct group in the context of medical image segmentation.
@@ 5.3 Ablation Study
5.3.1 Effects of APG and MLAM
The APG in our AutoProSAM refines the feature maps from the final stage of the attention block in the image encoder, paralleling the way the original SAM’s prompt encoder processes manual prompts. Without APG, these feature maps are directly fed into the mask decoder. Beyond APG, a Multi-Layer Aggregation Mechanism (MLAM) is also applied for the integration of information from various stages of the encoder and APG to the final mask decoder. To evaluate the effectiveness of both APG and MLAM, we conducted a comparative analysis using the BTCV dataset by removing these components. The findings, presented in Tab. 4, reveal significant improvements in both Dice and NSD metrics when APG and MLAM are employed. These results highlight the substantial contribution of APG and MLAM to enhancing the efficacy of medical image segmentation.
@@ 5.3.1 Effects of APG and MLAM
5.3.2 Inference Time Analysis
As demonstrated in Table 5, our AutoProSAM, despite incorporating a heavy ViT encoder, achieves an inference speed only marginally slower than that of nnUNet [11] and SwinUNETR [34]. Notably, it significantly outperforms both UNETR [9] and nnFormer [46] in terms of speed, highlighting its remarkable efficiency among these SOTA methods.
Model Inference Time (sec. per scan)
nnUNet [11] 3.74
nnFormer [46] 11.73
UNETR [9] 58.93
SwinUNETR [34] 4.24
AutoProSAM (ours) 7.06
Table 5: Averaged inference time for different models in BTCV experiments with a patch size of $128\times 128\times 128$.

### v4

@@ Abstract
Segment Anything Model (SAM) is one of the pioneering prompt-based foundation models for image segmentation and has been rapidly adopted for various medical imaging applications. However, in clinical settings, creating effective prompts is notably challenging and time-consuming, requiring the expertise of domain specialists such as physicians. This requirement significantly diminishes SAM’s primary advantage—its interactive capability with end users—in medical applications. Moreover, recent studies have indicated that SAM, originally designed for 2D natural images, performs suboptimally on 3D medical image segmentation tasks. This subpar performance is attributed to the domain gaps between natural and medical images and the disparities in spatial arrangements between 2D and 3D images, particularly in multi-organ segmentation applications. To overcome these challenges, we present a novel technique termed AutoProSAM. This method automates 3D multi-organ CT-based segmentation by leveraging SAM’s foundational model capabilities without relying on domain experts for prompts. The approach utilizes parameter-efficient adaptation techniques to adapt SAM for 3D medical imagery and incorporates an effective automatic prompt learning paradigm specific to this domain. By eliminating the need for manual prompts, it enhances SAM’s capabilities for 3D medical image segmentation and achieves state-of-the-art (SOTA) performance in CT-based multi-organ segmentation tasks. The code is in this link.
@@ 1 Introduction
Despite their extensive training, large foundation models such as SAM struggle to adapt to medical images. This is primarily due to the inherent differences between medical and natural images and the challenges associated with effective prompting (as shown in Figure 1). Recent efforts have attempted to extend the success of SAM to medical image segmentation tasks [46, 41, 23, 9, 35]. However, the demonstrated performance has exhibited reduced precision and stability, particularly in more intricate segmentation tasks in 3D medical images characterized by smaller sizes, irregular shapes, and lower contrast properties compared to 2D natural images [9]. Figure 1(A) illustrates this oversight of substantial domain disparities between natural and medical images. Furthermore, unlike natural images, medical images in various formats are not easily interpretable by general users. The prompts, intended to guide the foundation models to focus on a particular location, become a performance bottleneck as users struggle to identify the region of interest. Therefore, generating effective prompts for these models necessitates the involvement of domain experts and physicians. This reliance on labor-intensive manually generated prompts [7, 35] hampers its successful application, particularly in multi-organ medical image segmentation tasks, as demonstrated in Figure 1(B). As a result, these foundation models have been shown to underperform in medical image segmentation tasks compared to state-of-the-art (SOTA) models specifically designed for medical imagery [9].
@@
  • We conduct extensive experiments and analysis on three public and one private multi-organ segmentation dataset, demonstrating that the proposed AutoProSAM achieves superior performance in medical image segmentation tasks compared to the baselines.
@@ 2.2 Parameter-efficient Model Fine-Tuning
Given the extensive utilization of foundation models, the concept of parameter-efficient fine-tuning has garnered significant attention. Existing methods for efficient fine-tuning can be categorized into three groups [5]. First, addition-based methods incorporate lightweight adapters [29, 39] or prompts [22, 15] into the original model, focusing solely on adjusting these parameters. Second, specification-based methods [43, 10] concentrate on selecting a small subset of the original parameters for tuning. Third, reparameterization-based methods [12] leverage low-rank matrices to approximate parameter updates. Recently, some researchers have extended pre-trained image models to encompass video comprehension [29] or volumetric segmentation [39]. However, these methods treat the additional dimension as a “word group” and employ specialized modules to aggregate information along the word dimension. In contrast, our work considers all three dimensions simultaneously and directly adapts the trained transformer block to capture patterns in 3D medical image inputs.
@@ 3 Method
In this section, we explain how to modify the original SAM architecture, initially developed for 2D natural images, to work with 3D volumetric medical images for segmentation tasks. We begin by offering a brief overview of the SAM framework (as depicted in Figure 2), followed by a detailed explanation of the adjustments made to the image encoder, prompt encoder, and mask decoder.
@@ 3.2 Handling 3D Medical Inputs
To enhance SAM’s capability in processing 3D medical images, we propose an adaptation strategy named AutoProSAM, as illustrated in Figure 2A. This strategy has two primary objectives: firstly, to enable the model to directly learn 3D spatial patterns, and secondly, to ensure continuity by inheriting most parameters from the pre-trained model while introducing easily adjustable incremental parameters. The detailed design is elaborated as follows.
@@ 3.2.4 Bottleneck Modifications
By making the above adjustments, we can smoothly transition the 2D ViT into a 3D ViT, reusing most parameters. However, fully fine-tuning the 3D ViT can be resource-intensive. To address this, we propose using a lightweight adapter approach for efficient fine-tuning. The adapter comprises a down-projection linear layer and an up-projection linear layer, represented as $\operatorname{Adapter}(\mathbf{X})=\mathbf{X}+\operatorname{Act}(\mathbf{X}W_{\mathrm{Down}})W_{\mathrm{Up}}$. Here, $\mathbf{X}\in\mathbb{R}^{N\times C}$ is the original feature representation, $W_{\mathrm{Down}}\in\mathbb{R}^{C\times N^{\prime}}$ and $W_{\mathrm{Up}}\in\mathbb{R}^{N^{\prime}\times C}$ are down-projection and up-projection layers, and $\mathrm{Act}(\cdot)$ is the activation function. To enhance 3D spatial awareness, we include a depth-wise 3D convolution after the down-projection layer, as shown in Figure 2B. This enhancement improves the adapter’s utilization of 3D spatial cues.
@@ 3.3 Auto Prompt Generator
To overcome these limitations, we propose to use an Auto Prompt Generator instead of positional encoding to represent the prompt. The whole process is illustrated in Figure 2C. Instead of using manually generated points or bounding boxes, we directly take the output feature map after the last block of attention and depth adapter operation. This APG follows a fully convolutional neural (FCN) based encoder-decoder design that resembles 3D UNet [34]. This generator boasts a lightweight structure, leveraging 3D-based convolution operations, and can be effortlessly learned from scratch. This enables precise prompt generation tailored to different medical segmentation tasks. Notably, it eliminates the need for additional manually generated prompts, simplifying and expediting the multi-class medical image segmentation tasks.
@@ 3.4 Mask Decoder
The SAM mask decoder was originally designed to be lightweight, using stacks of 2D convolution layers. In our updated version, we replaced these with 3D convolutions for direct 3D mask generation. While the original design works well for natural images with large, distinct objects, volumetric medical image segmentation benefits from U-shaped networks with multi-level skip connections [13, 36]. This is crucial for medical images, where objects are smaller and have less distinct boundaries, requiring higher-resolution details for better discrimination.
To address this need while keeping the design lightweight, we incorporate a multi-layer aggregation mechanism (MLAM) [47, 9] in our decoder. We utilize the intermediate feature maps from stages 1-4 (as depicted in Figure 2A) of the image encoder, along with the prompt embedding from the APG, to enrich the mask feature map without compromising efficiency. For improved resolution detail, we upsample the mask feature map to match the original resolution, then concatenate it with the original image. This concatenated map is fused using another 3D convolution to generate the final mask. This method effectively combines high-resolution details with the original image data in the mask generation process. We have streamlined the original SAM by focusing solely on a targeted downstream task and omitting features like multi-task generation and ambiguity awareness. The backbone of the mask decoder primarily comprises lightweight 3D convolutional layers, which are known for their ease of optimization. Consequently, we train all decoder parameters from scratch.
@@ 4.1 Datasets
AMOS: We mainly focus on the CT parts of the AMOS dataset [14]. The publicly accessible AMOS-CT dataset [14] consists of 200 multi-contrast abdominal CT scans for training and 100 for testing. These scans are annotated for 15 anatomies, including spleen, right kidney, left kidney, gallbladder, esophagus, liver, stomach, aorta, inferior vena cava, pancreas, right adrenal gland, left adrenal gland, duodenum, bladder, and prostate/uterus.
@@ 4.2 Implementation Details
We provide detailed pre-processing information about the interpolation to a fixed voxel spacing, HU ranges, and intensity normalization in Table 1.
@@ 4.3 Loss Function
For training the AutoProSAM (as shown in Figure 2A), a combination of Dice loss and Cross-Entropy loss is used to assess the alignment between the predicted mask and the ground truth on a pixel-wise basis. The objective function for the segmentation head is defined as follows:
@@ 5 Results & Discussion
Model BTCV AMOS CT-ORG Institutional Pelvic
mDice $\uparrow$ mNSD $\uparrow$ mDice $\uparrow$ mNSD $\uparrow$ mDice $\uparrow$ mNSD $\uparrow$ mDice $\uparrow$ mNSD $\uparrow$
nnUNet [13] 84.34 73.21 87.43 77.12 85.51 75.60 88.04 83.87
AttUNet [38] 83.67 74.35 83.75 72.23 85.13 74.64 88.06 84.12
UNet++ [50] 79.33 72.84 84.32 73.34 84.87 74.21 87.53 82.65
nnFormer [49] 83.51 71.65 84.52 70.06 82.75 70.31 87.60 82.13
UNETR [11] 85.47 74.35 77.24 60.58 83.13 71.56 86.68 81.54
SwinUNETR [36] 86.58 75.26 86.19 74.83 84.34 72.32 89.12 83.54
Ours 87.15 78.83* 88.65* 79.41* 85.12 76.37* 91.30* 84.35*
@@ 5.1 Comparison with SOTAs
We extensively compare our model with the SOTA 3D medical image segmentation approaches, including the most recent Transformer-based methods including UNETR [11], SwinUNETR [36], and nnFormer [49], as well as CNN-based methods such as nnUNet [13], AttUNet [38], and UNet++ [50]. As indicated in Table 2, we observe that the proposed AutoProSAM generally outperforms other SOTA methods in both Dice and NSD metrics across all four datasets.
@@
The superiority of our AutoProSAM model is further highlighted by qualitative results. As depicted in Figure 3, qualitative comparisons of the predicted masks from various segmentation models reveal that our AutoProSAM yields visually superior mask predictions, particularly in terms of more accurate boundary delineation, when compared to its SOTA counterparts. In summary, while the original SAM displayed relatively weaker performance in medical image segmentation tasks as compared to SOTA methods [24], our adapted design registers significant enhancements and generalization to 3D medical image segmentation.
We also randomly selected samples from the CT-ORG dataset for more visual comparisons of different methods. As depicted in Figure 4, all methods generally achieve precise prediction masks for organs that are relatively easy to segment, such as lungs and bones. However, our AutoProSAM also provides more consistently accurate predictions than other baselines for organs like the liver, kidney, and bladder.
@@ 5.2 Comparison with SAM-based Methods
We further compare our method with existing SAM-based methods for multi-class medical image segmentation, including the original SAM and MedSAM [23]. MedSAM can be implemented with either full fine-tuning or partial fine-tuning, where only the prompt encoder and mask decoder are updated using point prompts. Other adaptation methods were not considered, as they either require collecting a large number of medical images for fine-tuning (e.g., Medical-SAM-Adapter [41], SAM-Med2D [2], SAM-Med3D [37]), or their models are too huge, where comparing would not be fair (Medical SAM Adapter [41] has tunable parameters approximately 336 million compared to our model of ten times fewer), or some models do not focus on organ segmentation (e.g., 3DSAM-Adapter [9] for tumor segmentation, MA-SAM [1] focuses on modality-agnostic performance for medical images), or some models are 2D models working with 2D slices (e.g., H-SAM [3] and SimTxtSeg [42]). Furthermore, we observe that many SAM-based works rely on ground truth images during the inference stage to generate prompts, which provides an unfair performance advantage. In contrast, our method, APG, does not depend on ground truth images, making it significantly more robust.
@@
The outcomes are meticulously outlined in Table 3, underscoring how our adaptation strategy surpasses all existing methods. Notably, our approach outshines the second-best technique by a margin of 3% in terms of BTCV segmentation Dice, 5% for NSD, and 6% concerning AMOS segmentation Dice, accompanied by a substantial 9% for NSD. Impressively, it even outperforms the complete fine-tuning variant of MedSAM, even when considering parameter-efficient fine-tuning. These results effectively validate our hypothesis that parameters pre-trained on 2D images can be effectively harnessed to grasp 3D spatial features with only minor adjustments. Moreover, our approach of treating all dimensions equivalently emerges as a superior strategy compared to interpreting the depth dimension as a distinct group in the context of medical image segmentation.
@@ 5.3 Ablation Study
Effects of APG and MLAM The APG in AutoProSAM refines feature maps from the final stage of the encoder, similar to how SAM’s prompt encoder processes manual prompts. Without APG, these maps go directly to the mask decoder. Additionally, the Multi-Layer Aggregation Mechanism (MLAM) integrates information from various stages into the final mask decoder. A comparative analysis on the BTCV dataset, shown in Table 4, demonstrates significant improvements in Dice and NSD metrics when APG and MLAM are used, highlighting their critical role in improving segmentation performance.
@@
Computational Requirements Analysis As demonstrated in Table 5, our AutoProSAM, despite incorporating a heavy ViT encoder, achieves an inference speed only marginally slower than that of nnUNet [13] and SwinUNETR [36]. Notably, it significantly outperforms both UNETR [11] and nnFormer [49] in terms of speed, highlighting its remarkable efficiency among these SOTA methods. Furthermore, AutoProSAM has the fewest tunable parameters during training, making it the most computationally feasible among the benchmark models.
Model Tuned Inference Time
Params. (M) (sec./scan)
nnUNet [13] 31.18 3.74
nnFormer [49] 150.14 11.73
UNETR [11] 93.02 58.93
SwinUNETR [36] 62.83 4.24
AutoProSAM (ours) 26.53* 7.06
Table 5: Averaged inference time for different models in BTCV experiments with a patch size of $128\times 128\times 128$.*Note: this may change slightly according to your exact setting of implementation.
Effectiveness for MRI Modality and More Organs To evaluate AutoProSAM, we compared its average performance on AMOS MRI [14] and MM-WHS CT (heart substructures) [51] datasets against benchmark models. The average dice scores, shown in Table 6, demonstrate AutoProSAM’s superiority across various configurations.
Model AMOS MRI MM-WHS CT
nnFormer 80.6 79.6
UNETR 75.3 85.8
SwinUNETR 75.7 86.1
AutoProSAM 80.8 90.9
Table 6: Comparisons on AMOS MRI [14] and MM-WHS CT [51] datasets on average Dice.